# A Note on the Phragmén-Lindelöf Theorem

### v2

@@ Abstract.
The Phragmén-Lindelöf principle is well known and well used in analytic number theory as a tool to obtain bounds on $L$-functions by interpolating bounds obtainable on vertical lines. One of the early general purpose results which is now widely used is [Rademacher, Theorem 2]. There is a generalization of this result presented in [Tr14-2, Lemma 3]. The proof of Rademacher is quite technical as it requires the construction of an analytic function with very precisely described growth properties. The proof of Trudgian has an error111the function $F(s)$ defined in the proof is not in general analytic as it is a function of $\sigma={\rm Re}(s)$. which arises in trying to avoid doing the same technical argument with power of $\log(Q+s)$ rather than $(Q+s)$.
In what follows we sufficiently correct the error so that proofs using it should be recoverable, at worst at the expense of some constants. Moreover, we provide a generalized version Rademacher’s work using what is in our view a simpler proof than that of Rademacher. We note that the result of Trudgian has been widely cited (45 direct citations, many of which have even more), including several results which now improve upon Trudgian’s main result (see [HSW, BW]). Consequently, that our result in practice recovers (at the expense of extra checks for small values, see Section 1) and improves upon Trudgian’s work (see Sections 2 and 3) means our work shall either recover these other results, or at least sets the groundwork for others to do so.
@@
Let $a,b$ be real numbers, and let $G_{1},\ldots,G_{r}$ be complex functions that are holomorphic for ${\rm Re}(s)\in[a,b]$, have $\overline{G(s)}=G(\overline{s})$, and satisfy the growth condition
@@
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$ for $\sigma\in[a,b]$. Suppose that $f$ is a holomorphic function on $a\leq\sigma\leq b$ and satisfies the growth condition
@@
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$. Suppose that we have real numbers $\alpha_{1},\ldots,\alpha_{r}\geq 0$ and $\beta_{1},\ldots,\beta_{r}\geq 0$ such that
$|f(a+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(a+it)^{\alpha_{i}}\bigg{|}$
@@
$|f(a+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(b+it)^{\beta_{i}}\bigg{|}.$
Finally, suppose $|G_{i}(\sigma+it)|$ is increasing in $\sigma$ if $\alpha_{i}>\beta_{i}$ and decreasing if $\beta_{i}<\alpha_{i}$. Then for $a\leq\sigma\leq b$, we have
$|f(\sigma+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(\sigma+it)^{\alpha_{i}}\bigg{|}^{\frac{b-\sigma}{b-a}}\bigg{|}\prod_{i=1}^{r}G_{i}(\sigma+it)^{\beta_{i}}\bigg{|}^{\frac{\sigma-a}{b-a}}.$
@@
The result can be extended to have $\alpha_{i}$ and $\beta_{i}$ negative at the expense of a growth condition
@@
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$ for $\sigma\in[a,b]$ by simply multiplying $f(s)$ by $G_{i}(s)^{-\min(\alpha_{i},\beta_{i})}$.
@@
Let $a$ and $b$ be real numbers and suppose $f(s)$ is holomorphic for ${\rm Re}(s)\in[a,b]$ and satisfies the growth condition
@@
then if $\sigma\in[a,b]$ we have $|f(\sigma+it)|$ is bounded by
@@
The function $|\log Q+s|$ is neither always increasing or always decreasing in $\sigma$. The behavior depends on $t$ and $\sigma$. For this reason the argument we gave above only works with $\alpha_{2}=\beta_{2}$ and $\alpha_{3}=\beta_{3}$.
Moreover, you should note well that $|\log(Q_{2}+s)|\geq\log|Q_{2}+s|$ and $|\log\log(Q_{2}+s)|\geq\log\log|Q_{2}+s|$. Indeed
@@
$|\log\log(Q_{2}+s)|=\bigg{(}1+\frac{{\rm arg}(\log(Q_{2}+s))^{2})}{(\log|\log(Q_{2}+s)|)^{2}}\bigg{)}^{1/2}\log|\log(Q_{2}+s)|$
@@
This difference can be relevant in both the hypothesis and conclusion of the theorem depending on the signs of the exponents. There are multiple results in the literature which fail to account for this. See Sections 2 and 3 for a method to improve this.
@@
We shall denote by $\overline{f}$ the analytic function $\overline{f}(s)=\overline{f(\overline{s})}$. We note that a bound on $|f(s)|$ in terms of $G_{i}(s)$ satisfying $\overline{G_{i}(s)}=G_{i}(\overline{s})$ implies the same bound on $|\overline{f}(s)|$.
@@
$H_{i}(s)=\begin{cases}G_{i}(a+s)^{\alpha_{i}}G_{i}(b-s)^{\beta_{i}}&\alpha_{i}>\beta_{i}\text{ and }G(a+\sigma+it)\text{ increasing for all }t\\ G_{i}(a+s)^{\beta_{i}}G_{i}(b-s)^{\alpha_{i}}&\alpha_{i}<\beta_{i}\text{ and }G(a+\sigma+it)\text{ decreasing for all }t\\ G_{i}(a+s)^{\alpha_{i}}G_{i}(b-s)^{\beta_{i}}&\alpha_{i}=\beta_{i}\end{cases}$
@@
$H(s)=\prod_{i=1}^{r}H_{i}(s)$
By our assumptions we have
@@
is analytic in $[0,(b-a)]$ and it satisfies the necessary growth condition to apply the standard Phragmén-Lindelöf principle (see for example [Rudin, Theorem 12.9]). Moreover, for ${\rm Re}(s)=0$ and ${\rm Re}(s)=(b-a)$ we have have $F(s)/H(s)\leq 1$ as
@@
Thus by the Phragmén-Lindelöf principle we have $F(s)/H(s)\leq 1$ on the whole interval. It follows by evaluating at $s=(b-a)/2+it$ that for all $t$ we have
$|F((a-b)/2+it)|\leq|H((a-b)/2+it)|=\prod_{i=1}^{r}|G_{i}((a+b)/2+it)|^{\alpha_{i}+\beta_{i}}.$
@@
where $\frac{n}{2^{k}}\in[0,1]$. The proof of this proceeds by induction on $k$, using Lemma 5 applied to the interval $a=\tfrac{\lfloor n/2\rfloor}{2^{k-1}}$ and $b=\tfrac{\lceil{n/2}\rceil}{2^{k-1}}$ with the bounds obtained from the inductive hypothesis.
@@
We note that in most applications as functions of $t$ we will have $|G_{i}(a+it)|\sim|G_{i}(\sigma+it)|\sim|G_{i}(b+it)|$ so that if one wants to work with a function where $G_{i}$ which is decreasing (respective increasing) in $\sigma$ rather then increasing (respective decreasing) one can replace $G_{i}(s)$ by $G(b-a-s)$ at the expense of possibly needing to multiply by a constant to preserve the inequalities on the boundary.
@@ 1. Dealing with the Condition of Increasing/Decreasing in $\sigma$
Even though the function $|\log Q+s|$ is neither always increasing or always decreasing in $\sigma$ it is desirable to be able to use it as $G_{i}(s)$. There are options for dealing with such functions
  • Replace $G(s)$ with $G(Q^{\prime}-s)$ provided the function was at least consistently always increasing or decreasing. This is essentially what occurs in [Rademacher, Theorem 2] for situations $\beta>\alpha$ applied to $G(s)=Q+s$.
@@
Let $a,b$ be real numbers, and let $G_{1},\ldots,G_{r}$ be complex functions that are holomorphic for ${\rm Re}(s)\in[a,b]$, have $\overline{G(s)}=G(\overline{s})$, and satisfy the growth condition
@@
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$ for $\sigma\in[a,b]$. Suppose that $f$ is a holomorphic function on $a\leq\sigma\leq b$ and satisfies the growth condition
@@
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$. Suppose that we have real numbers $\alpha_{1},\ldots,\alpha_{r}\geq 0$ and $\beta_{1},\ldots,\beta_{r}\geq 0$ such that
$|f(a+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(a+it)^{\alpha_{i}}\bigg{|}$
@@
$|f(a+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(b+it)^{\beta_{i}}\bigg{|}.$
Suppose that for $|t|>T_{0}$ that $|G_{i}(\sigma+it)|$ is increasing in $\sigma$ if $\alpha_{i}>\beta_{i}$ and decreasing if $\beta_{i}<\alpha_{i}$. Finally, suppose that for each $t\leq T_{0}$ we have
@@
$|f(\sigma+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(\sigma+it)^{\alpha_{i}}\bigg{|}^{\frac{b-\sigma}{b-a}}\bigg{|}\prod_{i=1}^{r}G_{i}(\sigma+it)^{\beta_{i}}\bigg{|}^{\frac{\sigma-a}{b-a}}.$
@@
The proof is as before, except, to obtain the that
@@
$|f(\sigma+it)|<\prod_{i}{\rm min}_{\sigma}(|G(\sigma+it)|)^{\alpha_{i}\frac{b-\sigma}{b-a}+\beta_{i}\frac{\sigma-a}{b-a}}$
one could check numerically on the whole region, alternatively, one can prove something stronger by using the maximum modulos condition to verify something on only the boundary.
@@ 2. Errors in the literature
There are unfortunately a number of errors in the literature where $|\log Q+s|$ was incorrectly handled as $\log|Q+s|$. So beyond the need to check the hypothesis which we needed in our corrected version the previous section would not allow one to fully recover the numerics of such results. In this section we highlight how one can potentially almost recover the numerics from such as error, at least for $t$ that are not too small.
@@
so that the hypothesis we have is actually stronger than a bound on $|\log\sigma+it|$. It would be desirable that the conclusions we prove would be similarly strong. As noted before the discrepancy here is that as $t\to\infty$
$|\log\sigma+it|=(1+O_{\sigma}(1/(\log|t|)^{2}))\log|t|.$
@@
$\tfrac{1}{2}\left|\log(1-(\sigma+it)^{2})\right|=(1+O_{\sigma}(1/|t|^{2})))\log|t|$
provides the core idea to improve this discrepancy. The essential idea here is that ${\rm arg}(1+\sigma+it)\sim-{\rm arg}(1-\sigma-it)$ so that we can obtain cancellation of error terms.
@@
Approximations for $\log$ and $\arctan$, which can be made precise, will give for $t>\!\!>0$ that
$\log\left(1+\left(\frac{Q_{1}+\sigma}{t}\right)^{2}\right)\left(1+\left(\frac{Q_{2}-\sigma}{t}\right)^{2}\right)\sim\left(\frac{Q_{1}+\sigma}{t}\right)^{2}+\left(\frac{Q_{2}-\sigma}{t}\right)^{2}$
@@
Then for $\sigma\in[a,b]$ we will have
$\log|t|<|G_{Q_{1},Q_{2}}(\sigma+it)|<(1+O_{Q_{1},Q_{2},\sigma}(1/t^{2}))\log|t|.$
Moreover, $|G(\sigma+it)|$ will be decreasing in $\sigma$ for $\sigma\in[-Q_{1},(Q_{2}-Q_{1})/2]$ and increasing for $\sigma\in[(Q_{2}-Q_{1})/2,Q_{2}]$ provided $t$ sufficiently large (depending on $Q_{1}$ and $Q_{2}$). Hence with a careful choice of $Q_{1}$ and $Q_{2}$ it can satisfy the conditions of Theorem 7.
@@
$|\log(G_{Q_{1},Q_{2}}(s))|=\left(1+O\left(\frac{1}{t^{2}(\log t)^{2}}\right)\right)\log|G_{Q_{1},Q_{2}}(s)|$
This has two applications. One, to introduce $\log\log$ terms we can use in Theorem 7, but also, that functions of the form
@@
so that one could potentially directly use it to prove a theorem like Theorem 7, again, potentially at the expense of introducing a constant.

### v3

@@ Abstract.
The Phragmén-Lindelöf principle is well known and widely used in analytic number theory as a tool to obtain bounds on $L$-functions by interpolating bounds obtainable on vertical lines. One of the early general purpose results which is now widely used is [undefg, Theorem 2]. There is a generalization of this result presented in [undefi, Lemma 3]. The proof of Rademacher is quite technical as it requires the construction of an analytic function with very precisely described growth properties. The proof of Trudgian contains an error111the function $F(s)$ defined in the proof is not in general analytic as it is a function of $\sigma={\rm Re}(s)$. which arises in trying to avoid doing the same technical argument with power of $\log(Q+s)$ rather than $(Q+s)$.
In what follows we sufficiently correct the error so that proofs using it should be recoverable, at worst at the expense of some constants. Moreover, we provide a generalized version Rademacher’s work using what is in our view a simpler proof than that of Rademacher. We note that the result of Trudgian has been widely cited (45 direct citations, many of which have even more), including several results which now improve upon Trudgian’s main result (see [undefb, undefa]). Consequently, that our result in practice recovers (at the expense of extra checks for small values, see Section 2) and improves upon Trudgian’s work (see Sections 3 and 4) means our work shall in the worst case allow one to almost recover these other result.
In Section 1 we present the simple form of our main theorem and its proof. In Section 2 we present a slightly more complicated version of the theorem which is more adaptable to many real world settings. In Section 3 we address a common error in the literature and illustrate how one can recover, within a very small margin, the values that would arise from making this error while simultaneously illustrating how to improve on what would normally have been the correct use of the theorem. Finally, in Section 4 we give several examples of applications of our theorems to the Riemann zeta function. In these examples we recover, within a factor of $(1+10^{-9})$ for $t>10^{5}$, and in many cases improve upon the erroneous applications we are aware of in the literature.
1. Simple form of Main Theorem
@@
Let $a$ and $b$ be real numbers and let $G_{1},\ldots,G_{r}$ be complex functions that are holomorphic for ${\rm Re}(s)\in[a,b]$, have $\overline{G_{i}(s)}=G_{i}(\overline{s})$, and satisfy the following growth condition for all $\sigma\in[a,b]$
@@
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$. Suppose that $f$ is a holomorphic function for ${\rm Re}(s)\in[a,b]$ and satisfies the growth condition
@@
for $\sigma\in[a,b]$. Suppose that we have real numbers $\alpha_{1},\ldots,\alpha_{r}\geq 0$ and $\beta_{1},\ldots,\beta_{r}\geq 0$ such that
$|f(a+it)|\leq\prod_{i=1}^{r}\left|G_{i}(a+it)\right|^{\alpha_{i}}$
@@
$|f(b+it)|\leq\prod_{i=1}^{r}\left|G_{i}(b+it)\right|^{\beta_{i}}.$
Finally, suppose that for $\sigma\in[a,b]$ and for all $t$ the function $|G_{i}(\sigma+it)|$ is increasing in $\sigma$ if $\alpha_{i}>\beta_{i}$ and decreasing if $\beta_{i}>\alpha_{i}$. Then for $a\leq\sigma\leq b$, we have
$|f(\sigma+it)|\leq\bigg(\prod_{i=1}^{r}\left|G_{i}(\sigma+it)\right|^{\alpha_{i}}\bigg)^{\frac{b-\sigma}{b-a}}\bigg(\prod_{i=1}^{r}\left|G_{i}(\sigma+it)\right|^{\beta_{i}}\bigg)^{\frac{\sigma-a}{b-a}}.$
@@
The result can be extended to have $\alpha_{i}$ and $\beta_{i}$ negative at the expense of adding the additional growth condition
@@
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$ for $\sigma\in[a,b]$ as a hypothesis. The conclusion is otherwise identical. The modified proof is to simply multiply $f(s)$ by $G_{i}(s)^{-\min(\alpha_{i},\beta_{i})}$.
@@
Let $a$ and $b$ be real numbers and suppose $f(s)$ is holomorphic for ${\rm Re}(s)\in[a,b]$ and satisfies the following growth condition for all $\sigma\in[a,b]$
@@
then if $\sigma\in[a,b]$ we have $|f(\sigma+it)|$ bounded by
@@
The function $|\log(Q+s)|$ is neither always increasing nor always decreasing in $\sigma$. The behavior depends on $t$ and $\sigma$. For this reason the argument we gave above only works with $\alpha_{2}=\beta_{2}$ and $\alpha_{3}=\beta_{3}$.
Moreover, you should recall that $|\log(Q_{2}+s)|\geq\log|Q_{2}+s|$ and $|\log\log(Q_{2}+s)|\geq\log\log|Q_{2}+s|$. Indeed
@@
$|\log\log(Q_{2}+s)|=\bigg(1+\frac{{\rm arg}(\log(Q_{2}+s))^{2}}{(\log|\log(Q_{2}+s)|)^{2}}\bigg)^{1/2}\log|\log(Q_{2}+s)|$
@@
This difference can be relevant in both the hypothesis and conclusion of the theorem depending on the signs of the exponents. There are multiple results in the literature which fail to account for this. See Sections 3 and 4 for a method to improve this.
@@
We shall denote by $\overline{f}$ the function $\overline{f}(s)=\overline{f(\overline{s})}$ which is analytic for ${\rm Re}(s)\in[a,b]$. We note that a bound on $|f(s)|$ in terms of $G_{i}(s)$ satisfying $\overline{G_{i}(s)}=G_{i}(\overline{s})$ implies the same bound on $|\overline{f}(s)|$.
@@
$H_{i}(s)=\begin{cases}G_{i}(a+s)^{\alpha_{i}}G_{i}(b-s)^{\beta_{i}}&\alpha_{i}>\beta_{i}\\ G_{i}(a+s)^{\beta_{i}}G_{i}(b-s)^{\alpha_{i}}&\alpha_{i}<\beta_{i}\\ G_{i}(a+s)^{\alpha_{i}}G_{i}(b-s)^{\beta_{i}}&\alpha_{i}=\beta_{i}\end{cases}$
@@
$H(s)=\prod_{i=1}^{r}H_{i}(s).$
The growth conditions in $G_{i}$ imply they are non-vanishing for ${\rm Re}(s)\in[a,b]$ which allows us to conclude that
@@
is analytic in [0,(ba)][0,(b-a)] and it satisfies the necessary growth condition to apply the standard Phragmén-Lindelöf principle (see for example [undefh, Theorem 12.9]). Moreover, for ${\rm Re}(s)=0$ and ${\rm Re}(s)=(b-a)$ we have $F(s)/H(s)\leq 1$ as
@@
Thus by the Phragmén-Lindelöf principle we have $F(s)/H(s)\leq 1$ on the whole interval. It follows by evaluating at s=(ba)/2+its=(b-a)/2+it that for all $t$ we have
$|F((b-a)/2+it)|\leq|H((b-a)/2+it)|=\prod_{i=1}^{r}|G_{i}((a+b)/2+it)|^{\alpha_{i}+\beta_{i}}.$
@@
where $n$ and $k$ are non-negative integers so that $\frac{n}{2^{k}}\in[0,1]$. The proof of this proceeds by induction on $k$, using Lemma 5 applied to the interval $a=\tfrac{\lfloor n/2\rfloor}{2^{k-1}}$ and $b=\tfrac{\lceil{n/2}\rceil}{2^{k-1}}$ with the bounds obtained from the inductive hypothesis.
@@
We note that in most applications as functions of $t$ we will have $|G_{i}(a+it)|\sim|G_{i}(\sigma+it)|\sim|G_{i}(b+it)|$ so that if one wants to work with a function where $G_{i}$ which is decreasing (respectively increasing) in $\sigma$ rather than increasing (respectively decreasing) one can replace $G_{i}(s)$ by G(bas)G(b-a-s) at the expense of possibly needing to multiply by a constant to preserve the inequalities on the boundary.
@@ 2. Dealing with the Condition of Increasing/Decreasing in $\sigma$
Even though the function $|\log(Q+s)|$ is neither always increasing nor always decreasing in $\sigma$ it is desirable to be able to use it as $G_{i}(s)$. There are options for dealing with such functions:
  • Replace $G(s)$ with $G(Q^{\prime}-s)$ provided the function was at least consistently always increasing or decreasing. This is essentially what occurs in [undefg, Theorem 2] for situations $\beta>\alpha$ applied to $G(s)=Q+s$. Notice that the function $|Q-\sigma-it|$ would be decreasing in $\sigma$ when $Q-\sigma>0$ whereas $|Q+\sigma+it|$ is increasing in $\sigma$ where $Q+\sigma>0$.
@@
Let $a,b$ be real numbers, and let $G_{1},\ldots,G_{r}$ be complex functions that are holomorphic for ${\rm Re}(s)\in[a,b]$, have $\overline{G_{i}(s)}=G_{i}(\overline{s})$, and satisfy the following growth condition for all $\sigma\in[a,b]$
@@
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$. Suppose that $f$ is a holomorphic function on ${\rm Re}(s)\in[a,b]$ and satisfies the growth condition
@@
for all $\sigma\in[a,b]$. Suppose that we have real numbers $\alpha_{1},\ldots,\alpha_{r}\geq 0$ and $\beta_{1},\ldots,\beta_{r}\geq 0$ such that
$|f(a+it)|\leq\prod_{i=1}^{r}\left|G_{i}(a+it)\right|^{\alpha_{i}}$
@@
$|f(b+it)|\leq\prod_{i=1}^{r}\left|G_{i}(b+it)\right|^{\beta_{i}}.$
Suppose for $|t|>T_{0}$ and $\sigma\in[a,b]$ that $|G_{i}(\sigma+it)|$ is increasing in $\sigma$ if $\alpha_{i}>\beta_{i}$ and decreasing if $\beta_{i}>\alpha_{i}$. Finally, suppose that for each $|t|\leq T_{0}$ we have
@@
$|f(\sigma+it)|\leq\bigg(\prod_{i=1}^{r}\left|G_{i}(\sigma+it)\right|^{\alpha_{i}}\bigg)^{\frac{b-\sigma}{b-a}}\bigg(\prod_{i=1}^{r}\left|G_{i}(\sigma+it)\right|^{\beta_{i}}\bigg)^{\frac{\sigma-a}{b-a}}.$
@@
The proof is as before, except that to obtain that
@@
$|f(\sigma+it)|<\prod_{i}{\rm min}_{\sigma}(|G_{i}(\sigma+it)|)^{\alpha_{i}\frac{b-\sigma}{b-a}+\beta_{i}\frac{\sigma-a}{b-a}}$
one could check numerically on the whole region using complex ball arithmetic, alternatively, one can prove something stronger by using the maximum modulus condition to verify something on only the boundary.
@@ 3. Errors in the literature
There are unfortunately a number of errors in the literature where $|\log(Q+s)|$ was incorrectly handled as $\log|Q+s|$. Consequently, correcting the literature in some cases would require more than simply checking the hypothesis needed in our corrected version; it would likely also require changes to the numerical results. In this section we highlight how one can almost recover the numerics from such an error, at least for $t$ that are not too small (for an illustration of what we mean by almost see Section 4).
@@
so that the input we have is actually stronger than the bound $|\log(\sigma+it)|$ which would normally be used as a hypothesis to our theorem. It would be desirable that the conclusions we prove would be similarly strong. As noted before the discrepancy here is that as $t\to\infty$
$|\log(\sigma+it)|=(1+O_{\sigma}(1/(\log|t|)^{2}))\log|t|.$
@@
$\tfrac{1}{2}\left|\log(1-(\sigma+it)^{2})\right|=(1+O_{\sigma}(1/|t|^{2}))\log|t|$
provides the core idea to improve this discrepancy. The essential idea here is that ${\rm arg}(1+\sigma+it)\sim-{\rm arg}(1-\sigma-it)$ so that we can obtain cancellation of error terms as in the formulas below.
@@
Approximations for $\log$ and $\arctan$, which can be made precise, will give as $t\to\infty$ that
$\log\left(\left(1+\left(\frac{Q_{1}+\sigma}{t}\right)^{2}\right)\left(1+\left(\frac{Q_{2}-\sigma}{t}\right)^{2}\right)\right)\sim\left(\frac{Q_{1}+\sigma}{t}\right)^{2}+\left(\frac{Q_{2}-\sigma}{t}\right)^{2}$
@@
Then for $\sigma\in[a,b]$ we will have, using $|1+i/t|=1+O(1/t^{2})$, that
$\log|t|<|G_{Q_{1},Q_{2}}(\sigma+it)|<\left(1+O_{Q_{1},Q_{2},\sigma}\left(\frac{1}{t^{2}\log|t|}\right)\right)\log|t|.$
Moreover, $|G_{Q_{1},Q_{2}}(\sigma+it)|$ will be decreasing in $\sigma$ for $\sigma\in[-Q_{1},(Q_{2}-Q_{1})/2]$ and increasing for $\sigma\in[(Q_{2}-Q_{1})/2,Q_{2}]$ provided $t$ sufficiently large (depending on $Q_{1}$ and $Q_{2}$). Hence with a careful choice of $Q_{1}$ and $Q_{2}$ it can satisfy the conditions of Theorem 7.
@@
$|\log(G_{Q_{1},Q_{2}}(\sigma+it))|=\left(1+O\left(\frac{1}{t^{2}(\log t)^{2}}\right)\right)\log|G_{Q_{1},Q_{2}}(\sigma+it)|$
provides a reasonable assymptotic approximation to $\log\log|t|$ One can use this directly to introduce $\log\log|t|$ terms while applying Theorem 7.
Another potentially useful application comes from observing that functions of the form
@@
so that one could potentially directly use it to prove a theorem like Theorem 7, at the expense of introducing a constant.